OpenStudy (anonymous):

Guys I need help with a problem. I have the first half figured out... Determine if triangle DEF with coordinates D (2, 1), E (3, 5), and F (6, 2) is an equilateral triangle. Use evidence to support your claim. If it is not an equilateral triangle, what changes can be made to make it equilateral? Be specific Well first off, We know that an equalaterial triangle has to be equal in length. SO I found the lengths of the sides. DE=17squared EF=18squared FD=17squared so at this moment its an isosceles not equal. When you move a point, it changes the length of two sides, not just one....

OpenStudy (anonymous):

using trig?

jimthompson5910 (jim_thompson5910):

Did you mean \[\Large DE = \sqrt{17}\] \[\Large EF = \sqrt{18}\] \[\Large FD = \sqrt{17}\] ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

But I need help

OpenStudy (anonymous):

sorry

jimthompson5910 (jim_thompson5910):

ok one second while I think over how to solve

jimthompson5910 (jim_thompson5910):

are you able to graph each point on the xy coordinate system?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what graphing tools do you have?

OpenStudy (anonymous):

microsoft paint

OpenStudy (anonymous):

I can make graphs

jimthompson5910 (jim_thompson5910):

have you heard of geogebra?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

I'm using that program and it's highly more accurate

jimthompson5910 (jim_thompson5910):

so go ahead and download and install that (it's a free program)

OpenStudy (anonymous):

ok I can use geogebra

OpenStudy (anonymous):

now what

jimthompson5910 (jim_thompson5910):

is the program running?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok do you see the "input bar" at the bottom?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

type in D = (2, 1), then hit enter, to plot the point D

jimthompson5910 (jim_thompson5910):

do the same for the other 2 points

jimthompson5910 (jim_thompson5910):

tell me when you have done that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

at the top, you should see this icon http://wiki.geogebra.org/uploads/1/1e/Tool_Circle_Center_Point.gif

jimthompson5910 (jim_thompson5910):

do you see it?

OpenStudy (anonymous):

so its equalaterial im not sure I can see it

jimthompson5910 (jim_thompson5910):

it looks equilateral, but you've proven above that it is not equilateral

jimthompson5910 (jim_thompson5910):

do you see that icon?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

in the corner

jimthompson5910 (jim_thompson5910):

click that icon. Then click on point D. And then click on point E You should see a circle pop up (centered at D and going through point E)

OpenStudy (anonymous):

Nothing is happening]

OpenStudy (anonymous):

can you explain to me what im supposed to be seeing

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

ok here are the three points D, E, F (see attached)

jimthompson5910 (jim_thompson5910):

Draw a circle that is centered at point D that goes through point E

jimthompson5910 (jim_thompson5910):

Notice how F is also on that circle. So this proves (at least visually) that DE = DF You've shown above that they are indeed equal because they both equal \(\Large \sqrt{17}\)

jimthompson5910 (jim_thompson5910):

Now draw a circle centered at point E and that goes through point D This circle will have the same radius as the last circle

OpenStudy (anonymous):

I can clearly see now that it is Isoseles

jimthompson5910 (jim_thompson5910):

notice how point F is NOT on the circle we just drew so this triangle is isosceles, but not equilateral

jimthompson5910 (jim_thompson5910):

to fix that, move point F to either point of intersection between the two circles

OpenStudy (anonymous):

now we have to make it equalaterial

jimthompson5910 (jim_thompson5910):

so move point F (in blue) to either point G (in red) or point H (in red)

jimthompson5910 (jim_thompson5910):

to analytically find those points G and H, you need to solve the system of equations (x - 2)^2 + (y - 1)^2 = 17 (x - 3)^2 + (y - 5)^2 = 17

OpenStudy (anonymous):

ok